What is the most important predictor of how much V̇O2max will decline with inactivity or bed rest? (make it understandable/right to the point if possible).
Answer #1

VO2 MAX- VO2 MAX is maximum amount of oxygen consumed in a minute to attain the physiological mechanism of lungs.

Factors which affect the VO2 MAX include- age, sex, type and mode of exercise practice, body composition and hereditary.

formula to calculate VO2 max ,

HRmax = calculate your maximum heart rate then subtract it from your age from 220.

If you’re 24 years old,

HRmax = 220 -24 = 196 beats per minute (bpm).

HRmin- check your heart rate while at rest (sitting down, doing little-to-no physical activity). The best time to measure your resting heart rate is in the morning before you get out of bed. place your two fingers over the radial artery andcount the ppulse for complete 60 s.

e. g - your heart rate in resting condition it is 72 b/minutes

VO2 max = 15 x (HRmax/HRrest).

The units for VO2 max are milliliters of oxygen per kilogram of body weight per minute (mL/kg/min).

from the above examples-  HRmax= 196beats per minute (bpm).,  HRmin= 72beats per minute (bpm).

then VO2 max = 15 x (HRmax/HRrest)

= 15 x (196/72)

= 15x 2.72

=40.8(mL/kg/min)

Non exercise variables i.e age, body mass and resting heart  can predict the VO2 MAX